# The Deligne–Simpson Problem

### v4

@@ 3. Connections on parabolic bundles
We will write $\Omega=\Omega^{1}_{\mathbb{P}^{1}}(\log D)$ for the sheaf of differential 1-forms on $\mathbb{P}^{1}$ with logarithmic poles on $D$. It is a line bundle on $\mathbb{P}^{1}$ of degree k2k-2, and for each point $a_{i}\in D$ there is a natural identification $\Omega_{a_{i}}\cong\mathbb{C}$. Explicitly, we can take as basis element $\frac{dz}{z}$ for any local co-ordinate around $a_{i}$ with $z(a_{i})=0$.
@@ 5. The extended tubular case
Suppose $\mathcal{F}=\mathcal{F}^{1}\oplus\mathcal{F}^{2}$ in $\operatorname{Coh}\mathbb{X}^{\prime}$. Then a fortiori $E=F^{1}\oplus F^{2}$ in $\operatorname{Coh}\mathbb{P}^{1}$, so the logarithmic connection $\nabla\colon E\to E\otimes\Omega$ decomposes as $(\nabla^{uv})$ for some $\mathbb{C}$-linear morphisms $\nabla^{uv}\colon F^{v}\to F^{u}\otimes\Omega$. The Leibniz identity tells us that $\nabla^{uu}$ is a logarithmic connection on $F^{u}$, whereas $\nabla^{uv}$ for $u\neq v$ is a morphism of sheaves. Similarly the residue $\operatorname{Res}_{a_{i}}\nabla$ decomposes as $(R_{i}^{uv})$, where $R_{i}^{uu}=\operatorname{Res}_{a_{i}}\nabla^{uu}$ is the residue of $\nabla^{uu}$, whereas $R_{i}^{uv}$ for $u\neq v$ is the fibre of the morphism $\nabla^{uv}$ (after canonically identifying the fibres of $F^{u}$ and $F^{u}\otimes\Omega$).
@@
Our basic strategy will be to deduce that $\phi^{uv}=0$ for some pair $u\neq v$. Lemma 2.2 will then tell us that $\nabla^{uv}$ is a morphism of parabolic sheaves $\mathcal{F}^{v}\to\mathcal{F}^{u}(\vec{\omega}_{\mathbb{X}^{\prime}})$ in $\operatorname{Coh}\mathbb{X}^{\prime}$. If we know that the only such homomorphism is zero, then we get an invariant parabolic subsheaf $\mathcal{E}^{v}$ of $(\mathcal{E},\nabla)$ by taking $\mathcal{F}^{v}$ and extending it by $E^{v}_{\infty}\coloneqq\operatorname{Im}\phi^{vv}$.
@@ 7. Multiplicative preprojective algebras
and a reflection functor $F_{q}$ from the category of modules for $\Lambda^{q}(Q)$ to modules for $\Lambda^{u_{v}(q)}(Q)$ acting as the reflection $s_{v}$ on dimension vectors.
Consider pairs $q,\alpha]$ with $q\in(K^{*})^{I}$ and $\alpha\in\mathbb{Z}^{I}$. The reflection at a loopfree vertex $v$ is admissible for the pair $q,\alpha]$ if $q_{v}\neq 1$. The equivalence relation $\sim$ on the set of pairs is generated by $q,\alpha]\sim[u_{v}(q),s_{v}(\alpha)]$ whenever the reflection at $v\in I$ is admissible for $q,\alpha]$. This is analogous to [1, §2], except that we use multiplicative reflections.
@@
This is an analogue of [1, Theorem 8.1], and is obtained by modifying the proof of that result, including straightforward analogues of [1, Lemmas 8.2-8.15] (8.8 and 8.13 need no change). For readers who want more details, we explain the steps.
For $\alpha\in\bar{F}_{q}$, we say that $\beta$ is a (1)(-1)-vector for the pair $q,\alpha]$ if $\beta,\alpha-\beta\in\mathbb{N}\bar{R}^{+}_{q}$ and $(\beta,\alpha-\beta)=-1$. We say that $\beta$ is a divisor for $q,\alpha]$ if it is a (1)(-1)-vector, $(\beta,\epsilon_{i})\leq 0$ for all $i$, and $(\alpha-\beta,\epsilon_{i})\leq 0$ whenever $(\beta,\epsilon_{i})=0$. If so, then
$1=(\beta,\alpha-\beta)=\sum_{i}(\alpha-\beta)_{i}(\beta,\epsilon_{i}),$
and the terms in the sum are $\leq 0$, so there is a unique vertex $j$, called the critical vertex for $\beta$, with $(\beta,\epsilon_{j})=-1$ and $(\alpha-\beta)_{j}=1$, and for every other vertex $i$ one has $(\beta,\epsilon_{i})=0$ or $(\alpha-\beta)_{i}=0$.
Let $\alpha\in\bar{F}_{q}\setminus\bar{\Sigma}_{q}$. Since $\alpha$ is an imaginary root, $q(\alpha)\leq 0$. Suppose first that $q(\alpha)=0$. By passing to an equivalent pair we may assume by the analogue of [1, Lemma 7.1], that $\alpha$ is in the fundamental region. Since $q(\alpha)=0$ this implies that the support of $\alpha$ is extended Dynkin and $\alpha$ is a multiple of the minimal positive imaginary root $\delta$. Let $m$ be the smallest positive integer with $(q^{\delta})^{m}=1$. Since $q^{\alpha}=1$, it follows that $\alpha=hm\delta$ for some $h\geq 1$. If $h=1$, then in any proper decomposition of $\alpha$ as a sum of roots $\alpha=\beta+\gamma+\cdots$ in $\bar{R}^{+}_{q}$, the terms must be real roots, and so $p(\alpha)>p(\beta)+p(\gamma)+\cdots$. Thus $\alpha\in\bar{\Sigma}_{q}$, a contradiction. We conclude that $h\geq 2$, so we are in case (I).
Suppose instead that $q(\alpha)<0$. Replacing $q,\alpha]$ by an equivalent pair we may ensure that $\alpha$ has support as small as possible. We then pass to the support quiver $Q^{\prime}$ of $\alpha$ and the restrictions $q^{\prime},\alpha^{\prime}]$ of $q$ and $\alpha$. Clearly $\alpha^{\prime}\in\bar{F}_{q^{\prime}}\setminus\bar{\Sigma}_{q^{\prime}}$. Observe that by minimality, if $q^{\prime\prime},\alpha^{\prime\prime}]$ is a pair equivalent to $q^{\prime},\alpha^{\prime}]$, then $\alpha^{\prime\prime}$ has support $Q^{\prime}$, and the pair $q^{\prime\prime},\alpha^{\prime\prime}]$ could be obtained from $q,\alpha]$ by first applying admissible reflections, and then passing to the support quiver of $\alpha^{\prime\prime}$.
By the analogue of [1, Lemma 8.3] there is a (1)(-1)-vector $\beta$ for $q^{\prime},\alpha^{\prime}]$, and hence a divisor $\beta^{\prime}$ for some equivalent pair $q^{\prime\prime},\alpha^{\prime\prime}]$ by the analogue of [1, Lemma 8.4]. Now $\beta^{\prime}$ and $\alpha^{\prime\prime}-\beta^{\prime}$ cannot both be sincere by the analogue of [1, Lemma 8.11]. Thus either $\beta^{\prime}$ is a non-sincere divisor, or we obtain a non-sincere divisor for some pair equivalent to $q^{\prime\prime},\alpha^{\prime\prime}]$ on applying the analogue of [1, Lemma 8.4] to the (1)(-1)-vector $\alpha^{\prime\prime}-\beta^{\prime}$ for $q^{\prime\prime},\alpha^{\prime\prime}]$. Thus case (II) or (III) holds by the analogues of [1, Lemmas 8.14 and 8.15]. ∎
@@
If there is a simple module, then by the reflection functors and Lemma 7.3, we may assume either that $\alpha$ is the coordinate vector of a loopfree vertex, or else $\alpha\in\bar{F}_{q}$. In the first case we have $\alpha\in\bar{\Sigma}_{q}$, so it is enough to consider $\alpha\in\bar{F}_{q}$. Again, because of the reflection functors of [7], it suffices to show that there is no simple module in the three cases (I), (II), (III) of Theorem  7.4.

### v5

@@ 2. Weighted projective lines and parabolic bundles
If $\mathcal{E}=(E,E_{ij})$ is a parabolic bundle and $F$ is a vector bundle, one gets a parabolic bundle $\mathcal{E}\otimes F$ by endowing the vector bundle $E\otimes F$ with the parabolic structure given by the subspaces
$E_{ij}\otimes F_{a_{i}}\subseteq E_{a_{i}}\otimes F_{a_{i}}\cong(E\otimes F)_{a_{i}}.$
Lenzing’s description of coherent sheaves on $\mathbb{X}$ as $p$-cycles of sheaves [14, §4.1] immediately gives the following.
Lemma 2.3.
We have $\mathcal{E}(\vec{c})\cong\mathcal{E}\otimes\mathcal{O}(1)$.
@@ 3. Connections on parabolic bundles
We will write $\Omega=\Omega^{1}_{\mathbb{P}^{1}}(\log D)$ for the sheaf of differential 1-forms on $\mathbb{P}^{1}$ with logarithmic poles on $D$. It is a line bundle on $\mathbb{P}^{1}$ of degree k2k-2 so isomorphic to $\mathcal{O}(k-2)$. For each point $a_{i}\in D$, the map sending a logarithmic form to its residue at $a_{i}$ induces a canonical isomorphism $\Omega_{a_{i}}\cong\mathbb{C}$. Explicitly, we can take as basis element $\frac{dz}{z}$ for any local co-ordinate around $a_{i}$ with $z(a_{i})=0$.
@@ 5. The extended tubular case
Suppose $\mathcal{F}=\mathcal{F}^{1}\oplus\mathcal{F}^{2}$ in $\operatorname{Coh}\mathbb{X}^{\prime}$. Then a fortiori $E=F^{1}\oplus F^{2}$ in $\operatorname{Coh}\mathbb{P}^{1}$, so the logarithmic connection $\nabla\colon E\to E\otimes\Omega$ decomposes as $(\nabla^{uv})$ for some $\mathbb{C}$-linear morphisms $\nabla^{uv}\colon F^{v}\to F^{u}\otimes\Omega$. The Leibniz identity tells us that $\nabla^{uu}$ is a logarithmic connection on $F^{u}$, whereas $\nabla^{uv}$ for $u\neq v$ is a morphism of sheaves. Similarly the residue $\operatorname{Res}_{a_{i}}\nabla$ decomposes as $(R_{i}^{uv})$, where $R_{i}^{uu}=\operatorname{Res}_{a_{i}}\nabla^{uu}$ is the residue of $\nabla^{uu}$, whereas $R_{i}^{uv}$ for $u\neq v$ is the composition
$F^{v}_{a_{i}}\to F^{u}_{a_{i}}\otimes\Omega_{a_{i}}\to F^{u}_{a_{i}}$
where the first map is the fibre of the morphism $\nabla^{uv}$ at $a_{i}$ and the second map is given by the canonical isomorphism $\Omega_{a_{i}}\cong\mathbb{C}$ mentioned at the start of §3.
@@
Our basic strategy will be to deduce that $\phi^{uv}=0$ for some pair $u\neq v$. Since $\vec{\omega}=(k-2)\vec{c}-\sum_{i=1}^{k}\vec{x}_{i}$, and $\Omega\cong\mathcal{O}(k-2)$, Lemmas 2.2 and 2.3 will then tell us that $\nabla^{uv}$ is a morphism of parabolic sheaves $\mathcal{F}^{v}\to\mathcal{F}^{u}(\vec{\omega}_{\mathbb{X}^{\prime}})$ in $\operatorname{Coh}\mathbb{X}^{\prime}$. If we know that the only such homomorphism is zero, then we get an invariant parabolic subsheaf $\mathcal{E}^{v}$ of $(\mathcal{E},\nabla)$ by taking $\mathcal{F}^{v}$ and extending it by $E^{v}_{\infty}\coloneqq\operatorname{Im}\phi^{vv}$.
@@ 7. Multiplicative preprojective algebras
and a reflection functor $F_{q}$ giving an equivalence from the category of modules for $\Lambda^{q}(Q)$ to modules for $\Lambda^{u_{v}(q)}(Q)$, and acting as the reflection $s_{v}$ on dimension vectors.
Consider pairs $q,\alpha]$ with $q\in(K^{*})^{I}$ and $\alpha\in\mathbb{Z}^{I}$. The reflection at a loopfree vertex $v$ is admissible for the pair $q,\alpha]$ if $q_{v}\neq 1$. The equivalence relation $\sim$ on the set of pairs is generated by $q,\alpha]\sim[u_{v}(q),s_{v}(\alpha)]$ whenever the reflection at $v\in I$ is admissible for $q,\alpha]$. This is analogous to [1, §2], except that we use multiplicative reflections. If $q,\alpha]\sim[q^{\prime},\alpha^{\prime}]$, then the reflection functors ensure that there is a simple $\Lambda^{q}(Q)$-module of dimension $\alpha$ if and only if there is a simple $\Lambda^{q^{\prime}}(Q)$-module of dimension $\alpha^{\prime}$.
@@
This is an analogue of [1, Theorem 8.1], and is obtained by modifying the proof of that result, including straightforward analogues of [1, Lemmas 8.2–8.15] (8.8 and 8.13 need no change). This result has been observed independently, and generalized, by Schedler and Tirelli [17, Theorem 6.16]. ∎
@@
If $q,\alpha]\sim[q^{\prime},\alpha^{\prime}]$, then by the multiplicative analogue of [1, Lemma 5.2(3)], we have $\alpha\in\bar{\Sigma}_{q}$ if and only if $\alpha^{\prime}\in\bar{\Sigma}_{q^{\prime}}$. Thus by Lemma 7.3, if there is a simple $\Lambda^{q}(Q)$-module of dimension $\alpha$, we may assume either that $\alpha$ is the coordinate vector of a loopfree vertex, or else $\alpha\in\bar{F}_{q}$. In the first case we trivially have $\alpha\in\bar{\Sigma}_{q}$, so it is enough to consider the case $\alpha\in\bar{F}_{q}$, and then it suffices to show that there is no simple module in the three cases (I), (II), (III) of Theorem  7.4.